# Deep Generative Models for Detector Signature Simulation: An Analytical Taxonomy

### v1

@@ 3.1.3 Representative Work and Datasets
Model
Algorithm
Representation
Conditioning
Experiment
Granularity $\uparrow$
IEA-GAN [45, 63]
GAN
grid/set
sensor position (radius and angle)
Belle II PXD (2023,2021)
$40\times 250\times 768=\mathbf{7,680,000}$ ch
WGAN [78]
GAN
grid
random
Belle II PXD (2019)
$40\times 250\times 768=\mathbf{7,680,000}$ ch
YonedaVAE [14]
VAE/ARM
multi-set
sensor position and Luminosity
Belle II PXD (2023)
$\mathbf{110,000}$ points
3DGAN  [61, 62]
GAN
grid
incident energy and angle
CLIC ECAL (2021, 2020)
$25\times 51\times 51=\mathbf{65,025}$ ch
BIB-AE [70]
VAE/GAN/NF
grid
incident energy and angle
ILD ECAL (2023)
$30\times 60\times 30=\mathbf{54,000}$ ch
CaloScore v2 [79]
Diffusion
grid
incident energy and time information
CaloChallenge D3 (2023)
$45\times 50\times 18=\mathbf{40,500}$ ch
iCaloFlow [80]
NF/ARM
grid
incident energy
CaloChallenge D3 (2023)
$45\times 50\times 18=\mathbf{40,500}$ ch
CaloDiffusion [81]
Diffusion
grid
incident energy
CaloChallenge D3 (2023)
$45\times 50\times 18=\mathbf{40,500}$ ch
CaloScore [82]
Diffusion
grid
incident energy and time information
CaloChallenge D3 (2022)
$45\times 50\times 18=\mathbf{40,500}$ ch
BIB-AE [83]
VAE/GAN
grid
incident energy
ILD AHCal (2021)
$48\times 25\times 25=\mathbf{30,000}$ ch
BIB-AE [84]
VAE/GAN
grid
incident energy
ILD ECAL (2021)
$30\times 30\times 30=\mathbf{27,000}$ ch
WGAN [85]
GAN
grid
energy and the deposit coordinates
EXO-200 (2023)
$74\times 350=\mathbf{25,900}$ ch
CaloClouds [86, 87]
VAE/Diffusion
multi-set
incident energy and cardinality
ILD ECAL (2023)
$\mathbf{6,000}$ points
NF-VAE [88]
VAE/NF
grid
random
Hyper-k IWCD (2019)
$19\times 16\times 40=\mathbf{12,160}$ ch
SuperCalo [89]
NF
grid
incident energy and geometry
CaloChallenge D2 (2023)
$45\times 16\times 9=\mathbf{6,480}$ ch
Variational Autoencoders. VAE is a simple yet flexible framework and could be adopted for controllable sampling by either modifying the loss function to enforce latent variables to be correlated with properties of interest or to feed the conditional information to different parts of the model. VAEs can also be incorporated to compress detector signatures in hybrid DGMs.
In [90], the ATLAS collaboration conditioned the encoder and decoder directly on the energy of the incident particle to generate showers corresponding to a specific energy. Abhishek et al. [88] utilized the VAE with the Normalizing Flows [31] to recover the true posterior distribution and to improve sampling for Water Cherenkov detector [91] simulation with grid-based data of size $19\times 16\times 40$.
In [92], the authors leverage the Sinkhorn Autoencoder [93] to have a trainable prior approximation, namely a noise generator to encode and generate embeddings with the same distribution in the latent space. In order to overcome the mode collapse issue and promote diversity, they include additional regularisation on the autoencoder’s latent space. Following [94], they compute a similarity matrix for the neural network’s weights according to the cosine similarity between its different layers to assess the diversity.
DeepRICH [95], designs a conditional latent space as a combination of CVAE [96] and infoVAE [97], where the latent variable $\sigma$ is determined using a Bayesian Optimization [98]. The control variables in this conditional VAE are the kinematic parameters of each particle learned by an auxiliary classifier over the encoded latent manifold as a regularization. They only consider the reconstruction of their dataset.
In [99, 84], the authors utilize the BIB-AE [100] model, conditioned on the incident photon energy. Thus, the latent manifold is conditioned on the Energy. They introduce a post-processing module that relaxes the trade-off between the accuracy of the emulated hit energy spectrum and the reproduced shower shape. This module is an MLP-based network that fixes the hit energy spectrum resolution between the input and generated images. In the later efforts [83] for hadronic showers, along with improving their model, inspired by [101], the authors use a Kernel Density Estimator (KDE) [102] to fit the learned latent manifold for the inference time and use rejection sampling for the correct density estimation. In an updated version of their work, McKeown et al. [70] add multi-parameter conditioning over the BIB-AE model, now conditioning on both incident angles and energies. Due to the multi-parameter space of their conditioning, they used Normalizing flows for latent space sampling during inference.
GVAE [103] introduces a graph-based VAE architecture for learning the representation of collision events without any controllable sampling for emulation. Orzari et al. [104, 105] develop a VAE for generating constituents of hadronic jets represented by sets. They incorporate a permutation invariant loss, Chamfer distance [106], instead of the typical mean squared error (MSE) as the reconstruction loss; however, their model does break the permutation equivariance using 2d convolution layers in the encoder. They regularize the model by constraining the $p_{T}$ and the invariant mass to follow the desired jet characteristics. The authors later in [107], fix this issue by using DGCNN [108] permutation equivariant layers.
Otten et al. [101] introduce a method called buffering density information given the encoded events. They construct a prior by aggregating a subset of the encoded training data by saving all the parameters of the Gaussian distributions for all events in the training data to a file, which constitutes the buffer. At inference time, to increase the variance and avoid overfitting to the training data, they also sample from the buffered Gaussian distributions with a variance control factor.
Collins et al. [109, 110], using a ParticleFlow [111] $\beta$-VAE [112], find an interpretable and meaningful representation of the jets and their information complexity by analyzing the VAE’s latent information. Moreover, they leverage $\beta$ from a fixed hyperparameter to an input of both the encoder and decoder networks.
Abhishek et al. [113] incorporated a Discrete Variational Autoencoder (DVAE) based model [114, 115, 116] with hierarchical dependencies of latent variables and a Restricted Boltzmann Machine (RBM) [117] latent prior using block Gibbs sampling for generation of the calorimeter showers. They also tackle the sparsity of the showers with a learnable masking like [118]. This solves the assignment problem with the sparsity, but it will be difficult to extend this to the simulation of high-granularity detector signatures.
Huang et al. [119], introduce an Autoencoder model for lossy compression of the data from Time Projection Chambers (TPCs) in sPHENIX experiment [120]. Their model, BCAE++, is an improvement over the original BCAE (Bicephalous Convolutional Autoencoder) [121] and is designed to handle the sparse data, achieving a higher compression ratio and better reconstruction accuracy. It utilizes a specialized loss function that combines a segmentation decoder for voxel-wise bi-class classification and a regression decoder for reconstruction. This dual-decoder approach allows the model to effectively handle the sparse and irregular distribution of the TPC data. A novel aspect of this work is the application of half-precision mode in the network without too much loss in reconstruction accuracy.
Cresswell et al. [122], develops a manifold hypothesis-inspired model (density estimation) [123, 124, 125] to make a dimensional reduction to the Calorimeter data to speedup the inference-level sampling process. This method comprises a two-step approach. Initially, the manifold of calorimeter showers is learned using a generalized Autoencoder, which aids in constructing low-dimensional latent encodings. Following this, density estimation is performed to capture a probability density within the learned manifold, alleviating the dimensionality mismatch typically encountered in maximum-likelihood estimation. This structured procedure allows for a more efficient and practical simulation of calorimeter showers. A possible drawback of this approach might be the presumption that the manifold encompasses showers, which could potentially limit the ability to generalize beyond the training data [126].
Hashemi et al. [14], marks the first model that robustly addresses the challenge of OOD simulation by introducing YonedaVAE, a novel multi-set generative model inspired by Category Theory [127, 128, 129]. YonedaVAE introduces a self-distilled set generator, capable of zero-shot creating sets and estimating the number of hits per sensor with variable “inter-event” and “intra-event” cardinality, facilitated by their Adaptive Top-q Sampling. Adaptive Top-q dynamically determines the multiplicity of points to be sampled based on the shape of the probability distribution for each event during inference. This adaptability is crucial for handling variable intra-event cardinality, especially when dealing with irregular detector geometries and hit patterns. They showed that YonedaVAE could produce new detector signature point clouds with cardinalities well beyond the training data and achieve context extrapolation. Trained on low luminosity data of PXD at Belle II with $\mathcal{O}(10^{2})$ hit multiplicity, YonedaVAE generates high luminosity valid signatures with the correct intra-event correlation and ultra-high granularity of $\mathcal{O}(10^{5})$ hits, without exposure to similar data during training.
Generative Adversarial Networks. GAN-based models, by design, allow easy implementation of controllable sampling due to introducing a sample discriminator given the desired properties, though taming its convergence is very difficult.
The application of GAN-based generative models as implicit density estimators was embarked on by [5] where they simulated 2D jet images for high energy W bosons and QCD jets as their conditional classes while introducing 2d locally connected layers (LAGAN). CaloGAN [68, 130, 131] employs the LAGAN layers to generate layer-wise two-dimensional images that were conditioned on the primary particle energy ranging uniformly from 1–100 GeV. Vallecorsa et al. [132, 133, 61] uses 3D ACGAN [134], to generate the calorimeter showers. In [62], they add incident angle conditioning as well. Erdmann et al. [135], uses WGAN [136] with continuous air shower energy conditioning using a constrainer network (like ACGAN). Musella et al. [118], for the generation of sparse hadronic jets using a U-net [137] module for the generator, introduce a decision-making method by adding an additional channel to their output as a mask probability to decide if a pixel should be zero or not. Srebre et al. [78] used the WGAN-gp [138] with random sampling to provide a proof of concept for the offline compression of the background hitmaps of ultra-high resolution Pixel Vertex Detector.
In [139], the authors do controlled sampling by conditioning the WGAN-gp model for the LHCb Calorimeter images. In [140], they introduce an emulator-simulator setup that benefits from the Siamese Network [141, 142, 143]. They show that using their parameter-to-image grid 2-stage training pipeline, they can model complex functions. In the pre-training stage, the goal is to learn an emulator distribution that matches the Monte-Carlo simulator distribution using the Siamese network to learn the similarity of the simulated and emulated images. Then, at the next stage, a generator will be trained to learn to map the random noise to the parameter space. All these stages follow an adversarial training regime.
Hashemi et al. [63], for the first time, generate ultra-high granularity PXD detector track hits with more than $7.5$ million pixel channels per event [74] — the highest spatial resolution detector simulation dataset ever analyzed with deep generative models. Building upon their previous work, the authors later introduce the Intra-Event Aware GAN (IEA-GAN) in [45]. As a fusion of the Transformer [144] model and GANs with self-supervised learning and conditional contrastive loss, they introduce the Relational Reasoning Module to approximate the concept of an “event” in sampling. They do the conditioning on the position of the PXD sensors (angles and radius) using methods from Deep Metric Learning [145]. Introducing self-supervised learning objectives in deep generative models, they show that they can generate fine-grained samples with large intra-event and small inter-event variations. Their evaluation encompasses a spectrum of results, extending from track-level (signature-level) features to the downstream performance of Helix (impact) parameter reconstruction.
Diefenbacher et al. [146] introduces a method for refining the precision of GANs using a post-processing re-weighting and tuning function based on [147, 148]. Kansal et al. [149, 150] for the first time choose a more sparse representation of the detector data and introduce a graph GAN based on Neural Message Passing layers [151]. Shirobokov et al. [152] introduce a new approach that synergizes deep generative models and non-differentiable simulators. They show that one can both approximate the stochastic behavior of the simulator and enable direct gradient-based optimization of an objective by parameterizing the latent variable model with the relevant parameters of the simulator. Jaruskova et al. [153], improves the calorimeter simulations over the lower energy depositions with AdaGAN-based [154] ensemble of GANs. Li et al. [85] use a WGAN model, conditioned on scintillation energy and the deposit coordinates, for simulating photodetector signals in the EXO-200 experiment’s time projection chamber with grid-shaped data of size $74\times 350$.
In [155], Winterhalder et al. introduce the Latent Space Refinement (LaSeR) protocol to enhance the precision and the topological obstruction of sampling by refining the predictions of a generator. In LaSeR, each generated sample is assigned a weight, which is then mapped to the corresponding latent space point. Rather than directly sampling from this weighted latent space, which could lead to biased results, the authors propose training a second generative model, the refiner, to transform the weighted latent space into an unweighted one.
CaloShowerGAN [156] employs Dataset 1 from the Fast Calorimeter Simulation Challenge 2022 [66] along with a well-structured training regime and data pre-processing to demonstrate revival of the classic CaloGAN [68]. This study defines a new GAN-based baseline and addresses the void in the baseline comparisons for calorimeter shower simulation, superseding the CaloGAN.
Hashemi et al. [157], for the first time, uses GAN to directly emulate high-level features computed from the reconstructed $Z\rightarrow\mu\mu$ events. DijetGAN [158] a simple GAN with random sampling for the simulation of QCD dijet events. Butter et al. [159], do top pair generation with a modified MMD-GAN [160] where the MMD kernel helps to describe on-shell resonances as well as tails of distributions, an improvement to [157]. They also study the statistical uncertainties and do ablation studies of the GAN approach to the event simulation using GANs. Carrazza et al. [161] employ CycleGAN [162] with the cycle-consistency loss to create mappings between two domains of Lund images [163], different categories of jets. Farrell et al. [164] apply GANs to generate full particle physics events conditioned on physics theory parameters.
Li et al. [165], introduce a style-based [166] conditional GAN [167] to predict lepton decay angles in the rest frames of $W$ bosons in Vector Boson Scattering (VBS) processes. Their approach addresses the challenge of missing neutrino information in the final state, which traditionally hampers the full determination of lepton angles. Alanazi et al. [168, 169] develop a GAN-based model that does an importance sampling over the generated features that improve the sensitivity of the discriminator. Prieto et al. [170] propose a style-based quantum GAN to generate events with a 3-qubit model. Howard et al. [171] incorporate the Sliced-Wasserstein VAE [172] and the theory-based physics constraints in an unsupervised setting for event generation. Epic-GAN [173], by Buhmann et al., unveils an equivariant point cloud module for simulating particle jets as multi-sets, addressing the computational challenges with existing models. The framework, grounded on deep sets, showcases significant computational efficiency by avoiding pairwise information sharing between jet constituents. A notable merit of EPiC-GAN lies in its scalability to large particle multiplicities for jet data.
Käch et al. in [174] propose a new GAN-based model that leverages attention mechanisms to generate particle cloud jets. The architecture is designed to handle a variable inter-event number of particles. They utilize a Bert-like [175] aggregation mechanism that scales linearly with the particle multiplicity. As a result, they introduce a “mean-field” particle, which serves as an aggregation point for information from individual particles in the cloud. This mean-field particle is initialized by the sum of all particles and is used in a cross-attention mechanism [144] to dynamically select which particles are important for distinguishing real from generated jets. This attention-based aggregation is particularly useful for handling large variances in individual particle energies. They achieve better results in jet simulation in comparison to the Epic-GAN [173] model.
Anderlini et al. [176] introduce a distilled GAN from an ensemble of models [177] to reduce the variance for a maximally diverse set of models. In the inference time, they test the performance of their model in the Out-Of-Distribution (OOD) regions of the phase space.
Flow-based Models. Flow-based Models [31, 32] learn an explicit bijection between the data distribution and a known distribution (usually a simple one, like a multivariate normal). This allows for exact likelihood computation and precise sampling but does not involve a compressed hidden latent space in the same way as models like VAEs or GANs.
For the first time, the CaloFlow [178] applies the normalizing flow to the simplified calorimeter geometry of CaloGAN [68] with $504$ cells. Their model is a combination of Masked Autoencoder for Distribution Estimation (MADE) blocks [179] and RQS transformations [180]. CaloFlow provides the additional benefit of tractable likelihoods with application to parameter inference for particle reconstruction. In CaloFlow II [181], they incorporate Knowledge distillation to transfer the probability density of the stronger model (teacher) to the student; an Inverse Autoregressive Flow (IAF) [182] model which is much faster in sampling. CaloFlow was then subsequently applied to CaloChallenge dataset 1 in [183].
JetFlow [184] employs Discrete Normalising Flows with multiple coupling layers. Each coupling layer divides input features into two sets, with one undergoing an identity transformation and the other a parameterized element-wise transformation. This configuration, with the addition of mass constraints, aids in modeling the generation of jets, showcasing a structured approach toward capturing more complicated correlations between the jet particles. Later, the authors [185] introduce a post-processing transformer encoder module, adversarially trained, to refine the output of the normalizing flow model. Xu et al. [186] develop a conditional Normalizing Flow based on [187] with the emphasis on modeling the correlation between the kinematic variables.
[…]

### v2

@@ 3.2 Representative Work and Datasets
Model Algorithm Representation Conditioning Experiment Granularity $\uparrow$
IEA-GAN [hashemi_ultra-high-resolution_2023, hashemi_pixel_2021] GAN grid/set sensor position (radius and angle) Belle II PXD (2023,2021) $40\times 250\times 768=\mathbf{7,680,000}$ ch
WGAN [srebre_generation_2020] GAN grid random Belle II PXD (2019) $40\times 250\times 768=\mathbf{7,680,000}$ ch
YonedaVAE [hashemi_deep_2024] VAE/ARM multi-set sensor position and Luminosity Belle II PXD (2023) $\mathbf{110,000}$ points
3DGAN  [khattak_fast_2021, belayneh_calorimetry_2020] GAN grid incident energy and angle CLIC ECAL (2021, 2020) $25\times 51\times 51=\mathbf{65,025}$ ch
BIB-AE [diefenbacher_new_2023] VAE/GAN/NF grid incident energy and angle ILD ECAL (2023) $30\times 60\times 30=\mathbf{54,000}$ ch
CaloScore v2 [mikuni_caloscore_2023] Diffusion grid incident energy and time information CaloChallenge D3 (2023) $45\times 50\times 18=\mathbf{40,500}$ ch
iCaloFlow [buckley_inductive_2023] NF/ARM grid incident energy CaloChallenge D3 (2023) $45\times 50\times 18=\mathbf{40,500}$ ch
CaloDiffusion [amram_denoising_2023] Diffusion grid incident energy CaloChallenge D3 (2023) $45\times 50\times 18=\mathbf{40,500}$ ch
CaloScore [mikuni_score-based_2022] Diffusion grid incident energy and time information CaloChallenge D3 (2022) $45\times 50\times 18=\mathbf{40,500}$ ch
BIB-AE [buhmann_hadrons_2021] VAE/GAN grid incident energy ILD AHCal (2021) $48\times 25\times 25=\mathbf{30,000}$ ch
BIB-AE [buhmann_getting_2021] VAE/GAN grid incident energy ILD ECAL (2021) $30\times 30\times 30=\mathbf{27,000}$ ch
WGAN [li_generative_2023] GAN grid energy and the deposit coordinates EXO-200 (2023) $74\times 350=\mathbf{25,900}$ ch
CaloClouds [buhmann_caloclouds_2023, buhmann_caloclouds_2023-1] VAE/Diffusion multi-set incident energy and cardinality ILD ECAL (2023) $\mathbf{6,000}$ points
NF-VAE [abhishek_variational_2019] VAE/NF grid random Hyper-k IWCD (2019) $19\times 16\times 40=\mathbf{12,160}$ ch
SuperCalo [pang_supercalo_2023] NF grid incident energy and geometry CaloChallenge D2 (2023) $45\times 16\times 9=\mathbf{6,480}$ ch
Variational Autoencoders. The work of Abhishek et al. [abhishek_variational_2019] is historically the first application of VAEs in detector simulation. They utilized the VAE with a Normalizing Flows [rezende_variational_2016] learnable prior, to recover the true posterior distribution better and to improve sampling for Water Cherenkov detector [proto-collaboration_hyper-kamiokande_2018] simulation with grid-based data of size $19\times 16\times 40$. This is one of the pioneering works in VAE applications for shower simulation that unfortunately got ignored entirely by the community.
In [ghosh_deep_2020], the ATLAS collaboration conditioned the encoder and decoder directly in a vanilla VAE, on the energy of the incident particle to generate showers corresponding to a specific energy.
In [deja_end--end_2020], the authors study the performance of the Sinkhorn Autoencoder [patrini_sinkhorn_2019] and leverage it to have a trainable prior approximation, namely a noise generator to encode and generate embeddings with the same distribution in the latent space. In order to overcome the mode collapse issue and promote diversity, they include additional regularisation on the autoencoder’s latent space. Following [ayinde_regularizing_2019], they compute a similarity matrix for the neural network’s weights according to the cosine similarity between its different layers to assess the diversity.
In the context of Cherenkov detectors, DeepRICH [fanelli_deeprich_2020], designs a conditional latent space as a combination of CVAE [sohn_learning_2015] and infoVAE [zhao_infovae_2018], where the latent variable $\sigma$ is determined using a Bayesian Optimization [snoek_practical_2012]. The control variables in this conditional VAE are the kinematic parameters of each particle learned by an auxiliary classifier over the encoded latent manifold as a regularization. They only consider the reconstruction of their dataset.
Moving to even higher-dimensional calorimeter setups, in [buhmann_decoding_2021, buhmann_getting_2021], the authors utilize the BIB-AE [voloshynovskiy_information_2019] model, conditioned on the incident photon energy. Thus, the latent manifold is conditioned on the Energy. They introduce a post-processing module that relaxes the trade-off between the accuracy of the emulated hit energy spectrum and the reproduced shower shape. This module is an MLP-based network that fixes the hit energy spectrum resolution between the input and generated images. In the later efforts [buhmann_hadrons_2021] for hadronic showers, along with improving their model, inspired by [otten_event_2021], the authors use a Kernel Density Estimator (KDE) [parzen_estimation_1962] to fit the learned latent manifold for the inference time and use rejection sampling for the correct density estimation. In an updated version of their work, McKeown et al. [diefenbacher_new_2023] add multi-parameter conditioning over the BIB-AE model, now conditioning on both incident angles and energies. Due to the multi-parameter space of their conditioning, they used Normalizing flows for latent space sampling during inference.
To address the sparsity of the datasets, GVAE [hariri_graph_2021] introduces a graph-based VAE architecture for learning the representation of collision events without any controllable sampling for emulation. Abhishek et al. [abhishek_calodvae_2022] incorporated a Discrete Variational Autoencoder (DVAE) based model [rolfe_discrete_2017, vahdat_dvae_2018, khoshaman_gumbolt_2019] with hierarchical dependencies of latent variables and a Restricted Boltzmann Machine (RBM) [montufar_restricted_2018] latent prior using block Gibbs sampling for generation of the calorimeter showers. They also tackle the sparsity of the showers with a learnable masking like [musella_fast_2018]. This solves the assignment problem with the sparsity, but it will be difficult to extend this to the simulation of high-granularity detector signatures.
Huang et al. [huang_fast_2023], introduce an Autoencoder model for lossy compression of the data from Time Projection Chambers (TPCs) in sPHENIX experiment [noauthor_bnl_nodate]. Their model, BCAE++, is an improvement over the original BCAE (Bicephalous Convolutional Autoencoder) [huang_efficient_2021] and is designed to handle the sparse data, achieving a higher compression ratio and better reconstruction accuracy. It utilizes a specialized loss function that combines a segmentation decoder for voxel-wise bi-class classification and a regression decoder for reconstruction. This dual-decoder approach allows the model to effectively handle the sparse and irregular distribution of the TPC data. A novel aspect of this work is the application of half-precision mode in the network without too much loss in reconstruction accuracy.
Cresswell et al. [cresswell_caloman_2022], develop a manifold hypothesis-inspired model (density estimation) [noauthor_representation_nodate, brehmer_flows_2020, brown_verifying_2023] to make a dimensional reduction to the calorimeter data to speedup the inference-level sampling process. This method comprises a two-step approach. Initially, the manifold of calorimeter showers is learned using a generalized Autoencoder, which aids in constructing low-dimensional latent encodings. Following this, density estimation is performed to capture a probability density within the learned manifold, alleviating the dimensionality mismatch typically encountered in maximum-likelihood estimation. This structured procedure allows for a more efficient and practical simulation of calorimeter showers. A possible drawback of this approach might be the presumption that the manifold encompasses showers, which could potentially limit the ability to generalize beyond the training data [balestriero_learning_2021].
To address the scalability issue of DGMs for ultra-high granularities with sparse representations, Hashemi et al. [hashemi_deep_2024], introduced YonedaVAE, a novel multi-set generative model inspired by Category Theory [eilenberg_general_nodate, de_haan_natural_2020, dudzik_graph_2022]. For the first time, they do an Out-Of-Distribution (OOD) detector simulation. Trained on low luminosity data of PXD at Belle II with $\mathcal{O}(10^{2})$ hit multiplicity, YonedaVAE generates high luminosity valid signatures with the correct intra-event correlation and ultra-high granularity of $110,000$ number of hits (cardinality), without exposure to similar data during training. YonedaVAE introduces a self-supervised set generator, capable of zero-shot creating sets and estimating the number of hits per sensor with variable “inter-event” and “intra-event” cardinality, facilitated by their Adaptive Top-q Sampling. Adaptive Top-q dynamically determines the multiplicity of points to be sampled based on the shape of the probability distribution for each event during inference. This adaptability is crucial for handling variable intra-event cardinality, especially when dealing with simulating the full detector with irregular detector geometries and hit patterns. They showed that YonedaVAE could produce new detector signature point clouds with cardinalities well beyond the training data and achieve context extrapolation. This work can also be categorized as an ARM as it simulates each detector sensor and layer with a causal Transformer [radford_improving_2018] with a proper positional embedding.
For event generation, Otten et al. [otten_event_2021] introduce a method called buffering density information given the encoded events. They construct a prior by aggregating a subset of the encoded training data by saving all the parameters of the Gaussian distributions for all events in the training data to a file, which constitutes the buffer. At inference time, to increase the variance and avoid overfitting to the training data, they also sample from the buffered Gaussian distributions with a variance control factor.
Orzari et al. [orzari_sparse_2021, touranakou_particle-based_2022] develop a VAE for generating constituents of hadronic jets represented by sets. They incorporate a permutation invariant loss, the Chamfer distance [fan_point_2016], instead of the typical mean squared error (MSE) as the reconstruction loss; however, their model does break the permutation equivariance using 2D convolution layers in the encoder. They regularize the model by constraining the $p_{T}$ and the invariant mass to follow the desired jet characteristics. The authors later in [tsan_particle_2021], fix this issue by using Dynamic Graph Convolutional (DGCNN) [wang_dynamic_2019] permutation equivariant layers.
Collins et al. [collins_exploration_2022, collins_machine-learning_2022], using a ParticleFlow [komiske_energy_2019] $\beta$-VAE [higgins_beta-vae_2022], find an interpretable and meaningful representation of the jets and their information complexity by analyzing the VAE’s latent information. Moreover, they leverage $\beta$ from a fixed hyperparameter to an input of both the encoder and decoder networks.
Generative Adversarial Networks.
The application of GAN-based generative models as implicit density estimators was embarked on by [de_oliveira_learning_2017] where they simulated 2D jet images for high energy W bosons and QCD jets as their conditional classes while introducing 2D locally connected layers (LAGAN). CaloGAN [paganini_calogan_2018, de_oliveira_controlling_2018, paganini_accelerating_2018] employs the LAGAN layers to generate layer-wise two-dimensional images that were conditioned on the primary particle energy ranging uniformly from 1–100 GeV. Vallecorsa et al. [khattak_three_2018, vallecorsa_3d_2019, khattak_fast_2021] uses 3D Auxiliary Classifier GAN (ACGAN) [odena_conditional_2017], to generate the calorimeter showers. In [belayneh_calorimetry_2020], they add incident angle conditioning as well. Erdmann et al. [erdmann_generating_2018], uses WGAN [arjovsky_wasserstein_2017] with continuous air shower energy conditioning using a constrainer network (like ACGAN). Musella et al. [musella_fast_2018], for the generation of sparse hadronic jets using a U-net [ronneberger_u-net_2015] module for the generator, introduce a decision-making method by adding an additional channel to their output as a mask probability to decide if a pixel should be zero or not. Srebre et al. [srebre_generation_2020] used the WGAN-gp [gulrajani_improved_2017] with random sampling to provide a proof of concept for the offline compression of the background hitmaps of ultra-high resolution Pixel Vertex Detector.
In [chekalina_generative_2019], the authors do controlled sampling by conditioning a WGAN-gp model for LHCb calorimeter images. In [alonso-monsalve_image-based_2020], they introduce an emulator-simulator setup that benefits from the Siamese Network [bromley_signature_nodate, chopra_learning_2005, koch_siamese_nodate]. They show that using their parameter-to-image grid 2-stage training pipeline, they can model complex functions. In the pre-training stage, the goal is to learn an emulator distribution that matches the Monte-Carlo simulator distribution using the Siamese network to learn the similarity of the simulated and emulated images. Then, at the next stage, a generator will be trained to learn to map the random noise to the parameter space. All these stages follow an adversarial training regime.
Hashemi et al. [hashemi_pixel_2021], for the first time, generate the full ultra-high granularity PXD detector track hits with more than $7.5$ million pixel channels per event [hashemi_ultra-high_nodate] — the highest spatial resolution detector simulation dataset ever analyzed with deep generative models in particle physics. Building upon their previous work, the authors later introduce the Intra-Event Aware GAN (IEA-GAN) in [hashemi_ultra-high-resolution_2023]. As a fusion of the Transformer [vaswani_attention_2017] model and GANs with self-supervised learning and conditional contrastive loss, they introduce the Relational Reasoning Module to approximate the concept of an “event” in full detector sampling. They do the conditioning on the position of the PXD sensors (angles and radius) using methods from Deep Metric Learning [kaya_deep_2019]. Introducing self-supervised learning objectives in deep generative models, they show that they can generate fine-grained samples with large intra-event and small inter-event variations. Their evaluation encompasses a spectrum of results, extending from track-level (signature-level) features to the downstream performance of Helix (impact) parameter reconstruction. It is noteworthy that calorimeter simulations typically focus on the simulation of particle showers from a single particle origin and in a small region of the calorimeter (i.e. specifying one detector component of $\mathbf{L}$), which could indeed capture some aspects of inter-layer correlation within the scope of a localized area. However, IEA-GAN’s approach extends this concept by considering the entire event with multiple-particle origins that encompass the full detector (i.e. in all detector components $\mathbf{L}$) as a whole, where correlations among different sensors (various angles and layers) become important within its readout window. Given the unique topology and geometry of the PXD as a highly granular tracking detector, this distinction is critical and allows for a more comprehensive simulation, capturing the complex interplay within an event across the entire detector rather than just the localized particle shower.
Diefenbacher et al. [diefenbacher_dctrgan_2020] introduce a method for refining the precision of GANs using a post-processing re-weighting and tuning function based on [andreassen_neural_2020, badiali_efficiency_2020]. Kansal et al. [kansal_graph_2021, kansal_particle_2022] for the first time choose a more sparse representation of the detector data and introduce a graph GAN based on Neural Message Passing layers [gilmer_neural_2017]. Shirobokov et al. [shirobokov_black-box_2020] introduce a new approach that synergizes deep generative models and non-differentiable simulators. They show that one can both approximate the stochastic behavior of the simulator and enable direct gradient-based optimization of an objective by parameterizing the latent variable model with the relevant parameters of the simulator. Jaruskova et al. [jaruskova_ensemble_2023], improve the calorimeter simulations over the lower energy depositions with AdaGAN-based [tolstikhin_adagan_2017] ensemble of GANs. Li et al. [li_generative_2023] use a WGAN model, conditioned on scintillation energy and the deposit coordinates, for simulating photodetector signals in the EXO-200 experiment’s time projection chamber with grid-shaped data of size $74\times 350$.
In [winterhalder_latent_2021], Winterhalder et al. introduce the Latent Space Refinement (LaSeR) protocol to enhance the precision and the topological obstruction of sampling by refining the predictions of a generator. In LaSeR, each generated sample is assigned a weight, which is then mapped to the corresponding latent space point. Rather than directly sampling from this weighted latent space, which could lead to biased results, the authors propose training a second generative model, the refiner, to transform the weighted latent space into an unweighted one.
CaloShowerGAN [giannelli_caloshowergan_2023] employs Dataset 1 from the Fast Calorimeter Simulation Challenge 2022 [michele_fast_2022] along with a well-structured training regime and data pre-processing to demonstrate revival of the classic CaloGAN [paganini_calogan_2018]. This study defines a new GAN-based baseline and addresses the void in the baseline comparisons for calorimeter shower simulation, superseding the CaloGAN.
For the event/jet simulation, Hashemi et al. [hashemi_lhc_2019] use a GAN to directly emulate high-level features computed from the reconstructed $Z\rightarrow\mu\mu$ events. DijetGAN [di_sipio_dijetgan_2019] uses a simple GAN with random sampling for the simulation of QCD dijet events. Butter et al. [butter_how_2019], do top pair generation with a modified MMD-GAN [li_mmd_2017] where the MMD kernel helps to describe on-shell resonances as well as tails of distributions, an improvement over [hashemi_lhc_2019]. They also study the statistical uncertainties and do ablation studies of the GAN approach to the event simulation using GANs. Carrazza et al. [carrazza_lund_2019] employ CycleGAN [zhu_unpaired_2020] with the cycle-consistency loss to create mappings between two domains of Lund images [dreyer_lund_2018], different categories of jets. Farrell et al. [farrell_next_2019] apply GANs to generate full particle physics events conditioned on physics theory parameters.
Li et al. [li_polarization_2022], introduce a style-based [karras_style-based_2019] conditional GAN [mirza_conditional_2014] to predict lepton decay angles in the rest frames of $W$ bosons in Vector Boson Scattering (VBS) processes. Their approach addresses the challenge of missing neutrino information in the final state, which traditionally hampers the full determination of lepton angles. Alanazi et al. [alanazi_simulation_2021, velasco_cfat-gan_2022] develop a GAN-based model that does an importance sampling over the generated features that improve the sensitivity of the discriminator. Prieto et al. [bravo-prieto_style-based_2022] propose a style-based quantum GAN to generate events with a 3-qubit model. Howard et al. [howard_learning_2022] incorporate the Sliced-Wasserstein VAE [kolouri_sliced-wasserstein_2018] and the theory-based physics constraints in an unsupervised setting for event generation. Epic-GAN [buhmann_epic-gan_2023], by Buhmann et al., unveils an equivariant point cloud module for simulating particle jets as multi-sets, addressing the computational challenges with existing models. The framework, grounded on deep sets, showcases significant computational efficiency by avoiding pairwise information sharing between jet constituents. A notable merit of EPiC-GAN lies in its scalability to large particle multiplicities for jet data.
Käch et al. in [kach_attention_2023] propose a new GAN-based model that leverages attention mechanisms to generate particle cloud jets. The architecture is designed to handle only a variable inter-event number of particles. They utilize a Bert-like [devlin_bert_2019] aggregation mechanism that scales linearly with the particle multiplicity. As a result, they introduce a “mean-field” particle, which serves as an aggregation point for information from individual particles in the cloud. This mean-field particle is initialized by the sum of all particles and is used in a cross-attention mechanism [vaswani_attention_2017] to dynamically select which particles are important for distinguishing real from generated jets. This attention-based aggregation is particularly useful for handling large variances in individual particle energies. They achieve better results in jet simulation in comparison to the Epic-GAN [buhmann_epic-gan_2023] model.
Anderlini et al. [anderlini_generative_2023] introduce a distilled GAN from an ensemble of models [malinin_ensemble_2019] to reduce the variance for a maximally diverse set of models. In the inference time, they test the performance of their model in the Out-Of-Distribution (OOD) regions of the phase space.
Flow-based Models.
[…]